OpenStudy (anonymous):

Evaluate. cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5

OpenStudy (anonymous):

\[\sqrt[3]{5}\sqrt{5}/\sqrt[3]{5}^5 \]

OpenStudy (anonymous):

a=3 just saying

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

it will be easier if you change roots into powers .ex. square root to 1/2, cube root to 1/3. Then, use the law of indices.

OpenStudy (anonymous):

law of indices?

OpenStudy (anonymous):

when multiplied, you ad powers, when divided, you sub powers

OpenStudy (midhun.madhu1987):

 try this please